### v1

@@ Abstract
Forest fire prediction involves estimating the likelihood of fire ignition or related risk levels in a specific area over a defined time period. With climate change intensifying fire behavior and frequency, accurate prediction has become one of the most pressing challenges in Artificial Intelligence (AI). Traditionally, fire ignition is approached as a binary classification task in the literature. However, this formulation oversimplifies the problem, especially from the perspective of end-users such as firefighters. In general, as is the case in France, firefighting units are organized by department, each with its terrain, climate conditions, and historical experience with fire events. Consequently, fire risk should be modeled in a way that is sensitive to local conditions and does not assume uniform risk across all regions. This paper proposes a new approach that tailors fire risk assessment to departmental contexts, offering more actionable and region-specific predictions for operational use. With this, we present the first national-scale AI benchmark for metropolitan France using state-of-the-art AI models on a relatively unexplored dataset. Finally, we offer a summary of important future works that should be taken into account. Supplementary materials are available on github.
@@ 1 Introduction
The immediate cost of wildfire suppression, compensation for damaged property, and the aftermath of lost agricultural productivity and tourism put a strain on local economies. For instance, the 2017 wildfires in the Mediterranean region resulted in estimated economic damages exceeding €200 million. From a human perspective, wildfires in France often lead to evacuations, sometimes involving thousands of residents and tourists. Health risks related to smoke inhalation and post-fire water contamination are also significant concerns [9, 3, 2]. Ecologically, these fires can cause irreversible harm to unique ecosystems. Additionally, wildfires contribute to topsoil erosion, threatening agriculture and leading to landslides in hilly terrains. Lastly, the emissions from these fires contribute not only to local air pollution but also to France’s national carbon footprint, further complicating efforts to combat climate change. According to reporterre, more than 4,000 hectares have burned in 2024.
2 State of the art
Constructing a wildfire prediction model has been central to wildfire research. We define the wildfire prediction in definition 1.
@@ 2 State of the art
One of the most widely used wildfire risk assessment tools is the Fire Weather Index (FWI), developed in Canada in 1969. This system generates several fire risk indicators based on meteorological inputs such as temperature, precipitation, and wind speed. In France, the FWI has been adapted and studied for local contexts. Pimont et al. [15] employed Generalized Additive Models (GAMs) to estimate fire occurrence in Bouches-du-Rhône using FWI and forest cover data. Dupire et al. [8] analyzed the vulnerability of forest ecosystems to fire in the French Alps. Among these studies, the Mediterranean basin emerges as the most extensively examined region, due to its heightened susceptibility to wildfires.
Several studies worldwide have analyzed monthly forest fire probabilities using various methods. Zhang et al.[17] analyzed monthly forest fire probabilities in Changsha, China, using a Random Forest model. They identified vegetation as a key factor for these predictions. Their dataset (2004–2021) employed a 1:1 sampling ratio of fire to non-fire points, which introduces bias by excluding many potential non-fire locations. Chen et al.[5] also employed a GNN to predict fire size in Portugal, using the publicly available MontesinhoFires dataset. While this dataset includes features such as the Fire Weather Index (FWI) and burn duration, it covers only 2000–2003 and is therefore outdated. Monthly fire prediction has also been studied in Australia by Cisneros et al.[6]. The paper adopts a two-stage deep graphical regression: it first predicts the monthly probability that each Australian SA1/SA2 region experiences at least one wildfire, using a graph-convolutional neural network logistic model. Conditional on an ignition, it then fits an extended generalised Pareto distribution to the square‑root burned area, so the same framework quantifies both ignition risk and the expected severity of the area burned.
Mutakabbir et al. [14] train straightforward deep neural-network classifiers and regressors on data generated by Spatio-Temporal Agnostic Sampling (STAS), which balances each wildfire with a non-fire sequence taken M months earlier from the exact same sensor location so that the model learns feature changes rather than absolute spatio-temporal cues. All reported metrics come from an 80%/20% random train‑test split of the 1998‑2018 records—no dedicated temporal hold‑out was used in the main experiments, so training and test samples are intermixed in time
Weekly fire prediction has been attempted by Michail et al.[12]. They used GraphCast (Lam et al.[11]) coupled with a temporal encoder achieving a solid global AUPRC of 0.64 for one-period-ahead prediction. However, when the same model is evaluated per region, its score in Europe collapses to only $0.20$—one of the weakest of all regions, underlining that the overall metric masks large spatial disparities. This gap shows that a single ”global” network does not automatically generalize across continents and argues for tailoring or fine-tuning models to specific geographic domains, a need the authors themselves acknowledge. Additionally, the original GraphCast architecture was not tested.
Reducing wildfire danger to a binary outcome masks the strong regional heterogeneity of risk in France. Two ignitions in low-exposure Brittany can signal a crisis, while the same count in Mediterranean departments is routine. Standard models flatten these contrasts, trivializing near-certain July fires in the south and exaggerating sparse events elsewhere. Such uniform treatment misguides resource allocation because danger must be expressed relative to local climatology and capacity. Additionally, we found no analysis of any French department other than the Mediterranean basin, whereas, as shown in Figure 1, other regions have a non-negligible number of fires. This was particularly true in 2022 as discussed by Hardy Karim [10]. We therefore present the first national-scale AI benchmark for metropolitan France, replacing binary labels with a region-aware multi-class scheme. Our framework maps risk at the departmental level, directly aligning predictions with operational firefighting jurisdictions. Using a relatively unexplored dataset, we evaluate several state-of-the-art models across four forecast horizons and two targets. While the prediction of the highest classes is still an issue, the comparison between multi-class and binary variants shows gains in predicting fires in the highest risk regions, along with a better target interpretation. We also provide a data-driven strategy for selecting the best model for each horizon. These findings open new directions, from dynamic class adaptation to region-specific covariate design, for truly nuanced wildfire-risk prediction.
Figure 1: French forest fire distribution across all departments between 2017-06-12 and 2023-01-01. For visualisation purposes, we only selected departments where the total number of forest fires was greater than 50.
@@ 3.1 Processing fire occurrence and burned area
BDIFF is an internet application designed to centralize all data on forest fires across French territory since 2006 and make this information available to the public and state services. Fire locations are referenced by the name of the city where they occurred; coordinates are not provided. Fire are defined, notably, by the burned area.
We collected fire data from June 12, 2017, to December 31, 2023. We generated daily features for each French department during this period. The years 2017–2020 and 2022 were used for training, 2021 for validation, and 2023 for testing. Due to the severity of fire activity in 2022, it was included in the training set, while a more typical year (2021) was selected for validation.
City names were used to spatially locate fire points and generate 2 $\mathrm{km}\times 2\mathrm{~{}km}$ rasters per department. Two targets were defined: number of fires (FO) and burned (BA) area, across 4 time ranges: current day and the next 7, 15, and 31 days. For each day and department, we summed each target to obtain the total number of fires and burned area that occurred. We transformed the classical binary and regression problem into a multi-class classification problem. For each department, we created an ordinal 5-class signal for both occurrence and burned area using the K-Means algorithm. Each sample with 0 fires was set to class 0, and positive samples were clustered into 4 classes, representing None, Normal, Medium, High, and Extreme fire occurrence or burned area. With this representation, we highlight typical occurrences and burned areas rather than absolute values. As the typical number of fire occurrences may vary by department, this representation seems most effective and interpretable for firefighters. Predictions are always made daily.
Figure 3 compares class distributions between the Mediterranean basin and the rest of France. While daily fire occurrence levels are similar nationwide (except for class 4), longer horizons reveal distinct seasonality and higher fire intensity in the Mediterranean. For the burned area, class 4 shows high variance, suggesting that extreme events are scattered and department-specific. This supports the idea that our clustering method enhances predictability but may obscure moderate events when extreme cases dominate a department’s historical profile. The resulting class distributions show a strong imbalance, with class 0 dominating. Class 1 is the most frequent among positive classes for both targets, especially for the burned area. A specific histogram for each horizon is available in the supplementary materials.
Figure 3: Distribution of fire occurrence (a) and burned area (b) classes risk in the Mediterranean basin and the rest of France. Histograms have been computed relative to the category -Mediterranean (orange) or not (blue)- taking all horizons together.
@@ 3.2 Processing Features
Table 1: Summary of features used in this study. ’-’ means same as above.
@@
Temperature 12h, 16h Meteostat Elevation Static IGN
Dew Point - - Forest landcover - -
Precipitation - - Landcover - Cosia
@@
Highway Static OSMNX [1] Past risk Daily Calculated
Population - Kontur Past risk burned area - -
@@
The features used in this study fall into four categories (Table 1): Meteorological, Topographic, Socio-Economic, and Historical. Most were rasterized into 2 km resolution datacubes for consistency, except for historical data, which were computed at the site level (see Figure 2).
- Meteorological data were calculated on an $11\times 1$ 11 grid per department. Fire indices followed standard methods and were reset annually. The precipitation index, based on Chen et al. [4], captures short-term rainfall variation and was computed directly on the 3D raster.
- Past risk and burned area features represent prior fire activity, processed using cubic kernel convolution (5 classes) and shifted by one day. The kernel size was based on the average fire sequence duration, with a 3-day inactivity threshold defining sequence boundaries.
@@
- Calendar features capture daily context (e.g., weekdays, holidays, curfews) and were encoded using CatBoost [16] based on nationwide fire totals. Aggregated statistics (mean, sum, min, max) were computed across encoders to represent calendar risk.
- Categorical features were also CatBoost-encoded using fire counts per pixel. Where needed, rasters were also split by subcategories (e.g., primary vs. secondary roads) to preserve detail lost at 2 km resolution. NDVI, NDSI, NDBI, NDMI, and NDWI (Normalized Difference Vegetation, Snow, Build-up, Moisture, and Water Index) are satellite indices calculated using Landsat 1 and 2 bands. These indices represent key environmental characteristics such as vegetation density, snow cover, built-up areas, soil moisture, and surface water presence—all of which are relevant for assessing fire risk. The COSIA database originally contains 18 land cover classes, but some classes (such as swimming pools) may not be particularly useful for forest fire analysis. Therefore, we applied a logical reduction to retain only the land cover classes most relevant to forest fires.
Except for Department, Historical, and Calendar data, all spatial features were aggregated (min, max, mean) by cluster. Finally, all variables—aside from Past risk and Past Burned Area—were standardized.
@@ 3.3 Features selection
This process produced a total of 261 features. Those with no variance and those highly correlated with each other (Pearson, Spearman, and Kendall correlation threshold of 0.95) were removed, retaining those with the highest variance. This selection resulted in 179 features. The final feature list is accessible in the supplementary materials. The final dataset is composed of 2,392 unique dates (with 1,944 dates corresponding to a fire in France), across 94 departments, for a total of 224,942 rows.
@@ 4 Models
Figure 4 shows the training and evaluation process, the models used, and the architectures of the deep learning models. A detailed description of the deep learning layer parameters and tree-based configuration is available in the supplementary materials. For each model independently, we address the data imbalance problem by testing various proportions of class 0, starting from $5\%$ to $100\%$, with a step of $5\%$. We select the percentage that maximizes the Intersection over Union on the validation set. Time series models (LSTM, GRU, DilatedCNN, and GRUGNN) were tested on three time series lengths (5, 10, 15). We selected the length based on the scores on the whole dataset. The selected lengths were tested for both targets at each horizon and are available in the supplementary materials. All models were trained with CrossEntropy loss. The learning rate for deep learning models was set to 0.00005. All model parameters were tuned based on validation performance, mainly on horizon 0. The FWI system is based on the classification proposed by Dupire et. al [7]. The objective of the GRUGNN is to test whether varying temporal features and static features should be included in the same process or not. We define spatial features as those that are static (e.g., topography), while temporal features vary daily (e.g., meteorology)
Figure 4: Training and evaluation process and models architecture used in this article. Evaluation part show a comparison between FWI, LSTM, and Catboost at day 0. A full resolution image is available in supplementary materials.
@@
    Intersection over Union (IoU), which measures how well the predicted risk aligns with actual risk when an event occurs. IoU is well-suited for multiclass wildfire prediction as it accounts for class uncertainty and preserves class ordinality—predicting class 1 instead of 4 is penalized less than predicting 0.

### v2

@@ Abstract
Forest fire prediction involves estimating the likelihood of fire ignition or related risk levels in a specific area over a defined time period. With climate change intensifying fire behavior and frequency, accurate prediction has become one of the most pressing challenges in Artificial Intelligence (AI). Traditionally, fire ignition is approached as a binary classification task in the literature. However, this formulation oversimplifies the problem, especially from the perspective of end-users such as firefighters. In general, as is the case in France, firefighting units are organized by department, each with its own terrain, climate conditions, and historical experience with fire events. Consequently, fire risk should be modeled in a way that is sensitive to local conditions and does not assume uniform risk across all regions. This knowledge is generally ignore in the current researches on wildfire prediction with AI. This paper proposes a new approach and a new database that tailors fire risk assessment to departmental contexts, offering more actionable and region-specific predictions for operational use. With this, we present the first national-scale AI benchmark for metropolitan France using state-of-the-art AI models on a relatively unexplored dataset. Finally, we offer a summary of important future works that should be taken into account. Supplementary materials are available on github NicolasCaronPro/Localized-Forest-Fire-Risk-Prediction-ADepartment-Aware-Approach-for-Operational-DecisionSupport.
@@ 1 Introduction
The immediate cost of wildfire suppression, compensation for damaged property, and the aftermath of lost agricultural productivity and tourism put a strain on local economies. For instance, the 2017 wildfires in the Mediterranean region resulted in estimated economic damages exceeding €200 million, according to Le Monde (2017). From a human perspective, wildfires in France often lead to evacuations, sometimes involving thousands of residents and tourists. Health risks related to smoke inhalation and post-fire water contamination are also significant concerns [1, 2, 3]. Ecologically, these fires can cause irreversible harm to unique ecosystems. Additionally, wildfires contribute to topsoil erosion, threatening agriculture and leading to landslides in hilly terrains. Lastly, the emissions from these fires contribute not only to local air pollution but also to France’s national carbon footprint, further complicating efforts to combat climate change. According to Reporterre (2024), more than 4,000 hectares have burned in 2024.
1.1 Problem formulation
In this work, the problem of predicting the future forest fire risk is considered. We define the wildfire prediction in definition 1.
@@ 1.1 Problem formulation
1.2 State of the art
In the literature, we found publicly available datasets used to address the wildfire prediction problem.
Kondylatos et al. [4] released a dataset for wildfire hazard prediction in Greece, covering the period 2009–2021 at a daily spatial resolution of approximately $1,\mathrm{km}^{2}$. The dataset integrates meteorological variables, land cover, and population data. Each file is about 23 GB in size, which raises significant challenges for memory usage and accessibility.
The SeaFire Cube platform [5] offers 21 years of global data (2001–2021), 8-day intervals, $0.25^{\circ}$ resolution. It includes atmospheric, climate, vegetation, socioeconomic, and fire variables (burned area, CO2). Easy to access with tutorials, it suits seasonal prediction but not fine-scale management. Michail et al. [6] used it with GraphCast [7], reaching global AUPRC $0.64$, but only $0.20$ in Europe, highlighting poor regional generalization and the need for domain-specific models.
The EO4WildFires benchmark [8] targets severity forecasting (affected area). It combines multi-sensor series (Sentinel-1, Sentinel-2, NASA Power) for 31,730 events across 45 countries (2018–2022), annotated with EFFIS ($\sim$25 GB in one file). It is rich in features (temperature, precipitation, soil moisture, snow), but large volume hampers usability.
Mesogeos [9] provides a dataset for the Mediterranean region spanning 2006–2022, with a daily spatial resolution of approximately $1,\mathrm{km}^{2}$ and 27 variables covering meteorology, vegetation, land cover, and human activity. It includes records of ignitions and burned areas larger than 30 ha. Despite the availability of extraction tools and a leaderboard, the dataset’s large volume poses significant challenges. Moreover, excluding smaller fires also removes important signals, such as detection time and response efficiency. Among the evaluated models (LSTM, GTN, and Transformer), LSTM achieved the best performance.
Across all these datasets, the wildfire risk prediction problem is typically treated as a binary classification task (using the predicted probability to generate risk maps). Employing a very fine spatial resolution (1 km $\times$ 1 km) allows for a detailed representation of spatial variables (e.g., land cover, vegetation dryness). However, this also introduces a strong bias in the evaluation of metrics, since only a tiny fraction of the test samples correspond to fire events (e.g., only 1,228 fire pixels in [4] out of several million pixels). Furthermore, the aspect of model calibration has rarely been studied and models often replicate known high-risk zones, while it has been proved that accuracy drops at fine scales [10, 11]. Another limitation is the large memory consumption required, which restricts their use in practice. Finally, one element not provided by current datasets is information on the organization of firefighting units within each country, even though risk prediction should ideally be aligned with this scale of organization.
1.3 Contribution
Our contribution rests on two main points:
  1. 1.
    Unlike existing public datasets, ours integrates the administrative structure of French departments, which aligns with the fire service organization and makes predictions more realistic for operational use. We incorporate calendar features (day, holidays, etc.), detailed forest cover information (tree species), refined land cover, and a comprehensive set of fire indices (FWI, Nesterov, Angström). Spatial data are provided in two formats to facilitate categorical handling. All features are aggregated into an $2,\mathrm{km}\times 2,\mathrm{km}\times 1$ day xarray. To manage memory, only one departmental cube is loaded at a time—this increases aggregation time but enables flexible study areas ranging from a single department to the whole of France. We demonstrate usage through fire count prediction with an ordinal classification scheme, though the dataset also supports broader applications in risk management and analysis. Similar studies to [4, 9] can likewise be conducted.
  2. 2.
    Binary wildfire danger ignores France’s regional heterogeneity: two ignitions in low-risk Brittany may indicate a crisis, while the same in Mediterranean areas is routine. Standard models flatten these contrasts, trivializing near-certain summer fires in the south and exaggerating sparse events elsewhere. This uniformity misguides resource allocation. The prediction of the number of fires (or of the total burned area) is very random (4 vs. 5 vs. 6 fires), which may also limit the model’s convergence. We thus propose the first national AI benchmark for metropolitan France, replacing binary labels with a region-aware multi-class scheme.
@@ 2 Constructing the database
In this section, we present the original fire source we used to construct the database along with the original features sources.
@@ 2.1 Processing Fire occurrence and Burned Area
BDIFF is an internet application designed to centralize all data on forest fires across French territory since 2006 and make this information available to the public and state services. Fire locations are referenced by the name of the city where they occurred; coordinates are not provided. Fire are defined, notably, by the BA.
Data from June 12, 2017, to December 31, 2024 were collected. The years 2017–2020 and 2022 were used for training, 2021 and 2024 for validation, and 2023 for testing. Due to the severity of fire activity in 2022, it was included in the training set, while a more typical years was selected for validation. We did not include 2024 in the test split but used it for validation, because it was a relatively low-fire year (about 1,600 fires vs. 2,300 in 2023), which would make it a less reliable test year.
City names were used to geolocate fire points and to generate $2,\mathrm{km}\times 2,\mathrm{km}$ rasters for each department. Two daily targets were defined: Fire Occurrence (FO) and Burned Area (BA). Instead of treating these as binary or regression tasks, we reformulated them as multi-class problems by constructing an ordinal five-class signal for both FO and BA using K-Means clustering. Class 0 corresponds to the absence of fire, while positive samples were grouped into four levels (Normal, Medium, High, Extreme). This approach emphasizes typical rather than absolute values, thereby adapting to departmental variability and improving interpretability for firefighting operations. All predictions are made at the daily scale.
Figure 2 compares class distributions between the Mediterranean basin and the rest of France. Daily FO levels are similar nationwide (except for class 3 and 4, which are concentrated in the Mediterranean). For the BA, class 4 shows high variance, suggesting that extreme events are scattered and department-specific. This supports the idea that our clustering method enhances predictability but may obscure moderate events when extreme cases dominate a department’s historical profile. The resulting class distributions show a strong imbalance, with class 0 dominating. Class 1 is the most frequent among positive classes for both targets. Superior classes (2,3 and 4) are underrepresented, specially in BA.
Figure 2: Distribution of FO (a) and BA (b) classes risk in the Mediterranean basin and the rest of France. Histograms have been computed relative to the category -Mediterranean (orange) or not (blue)- taking all horizons together.
@@ 2.2 Processing Features
Table 1: Summary of features used in this study. ’-’ means the same as above.
@@
Temperature 12h, 16h Meteostat Elevation Static CourbesDeNiveau (IGN)
Dew Point - - Forest landcover - BDForet (IGN)
Precipitation - - Landcover - Corine
@@
Highway Static BDRoute (IGN) Past risk Daily Calculated
Population - Kontur Past risk BA - -
@@
Department - -
The features used in this database fall into four categories (Table 1): Meteorological, Topographic, Socio-Economic, and Historical. Most were rasterized into 2 km resolution datacubes for consistency, except for historical data, which were computed at the site level (see Figure 1).
- Meteorological data were calculated on an $11\times 11$ grid per department. Fire indices followed standard methods and were reset annually. The precipitation index, based on Chen et al. [12], captures short-term rainfall variation and was computed directly on the 3D raster.
- Past risk and BA features represent prior fire activity, processed using cubic kernel convolution (5 classes) and shifted by one day. The kernel size was based on the average fire sequence duration, with a 3-day inactivity threshold defining sequence boundaries.
@@
- Calendar features capture daily context (e.g., weekdays, holidays, curfews) and were encoded using CatBoost [14] based on departmental fire totals. Aggregated statistics (mean, sum, min, max) were computed across encoders to represent calendar risk.
- Categorical features were also CatBoost-encoded using fire counts per pixel. Where needed, rasters were also split by subcategories (e.g., primary vs. secondary roads) to preserve detail lost at 2 km resolution. NDVI, NDSI, NDBI, NDMI, and NDWI (Normalized Difference Vegetation, Snow, Build-up, Moisture, and Water Index) are satellite indices calculated using Landsat 1 and 2 bands. These indices represent key environmental characteristics such as vegetation density, snow cover, built-up areas, soil moisture, and surface water presence—all of which are relevant for assessing fire risk. The Corine database originally contains 44 land cover classes, but some classes were merged in order to aggregate similar information (such as snow and rock simulate no vegetation area). Therefore, logical reduction was applied to retain only the land cover classes most relevant to forest fires: Urban, transport, forest, natural vegetation, agriculture, grassland, natural non-vegetation area, littoral, water area, and wetland. Code is available in the supplementary materials.
Except for Department, Historical, and Calendar data, all spatial features were aggregated (min, max, mean) by cluster. Finally, all variables—aside from Past risk and Past BA—were standardized.
Storing all x-array files requires about 150 GB, with each data cube averaging  1.5 GB. Building all x-arrays can take several days - up to a week. We used a Dell Precision 7780 with 32 GB of RAM and a 13th Gen Intel® Core™ i7-13850HX (28 cores). We do not currently have the rights to publish the xarray files, but the code for downloading and reconstructing them is released on GitHub. Automated downloading of raw data is provided where possible; however, some features (e.g., CORINE and GEE) require authentication. All steps is documented.
@@ 2.3 Features selection
This process produced a total of 261 features. Those with no variance and those highly correlated with each other (Pearson, Spearman, and Kendall correlation threshold of 0.95) were removed, retaining those with the highest variance. This selection resulted in 162 features. The final feature list is accessible in the supplementary materials. The final dataset is composed of 2,758 unique dates (with 2,192 dates corresponding to a fire in France), across 92 departments (Paris was not include due to lack reliability concerning fire. Corse was not include due to the different vegetation compared to the rest of France), for a total of 253,736 rows.
@@ 3 Models
The models used are described in detail in table 2, description of the deep learning layer parameters, and tree-based configuration is available in the supplementary materials. The data imbalance problem was addressed by testing various proportions of class 0, starting from $5\%$ to $100\%$, with a step of $5\%$. The percentage selected maximize the Intersection over Union on the validation set. Time series models (LSTM, GRU, DilatedCNN) were tested with 10 days sequences. All models were trained with CrossEntropy loss. The learning rate for deep learning models was set to 0.00005. All model parameters were tuned based on validation performance, mainly on horizon 0. The FWI system is based on the classification proposed by Dupire et. al [15]. Standard deviation has been calculating at each step of the training process with 5 runs.
Table 2: Models used for the benchmark.
Model Description
Logistic Regression A linear model that estimates class probabilities using the logistic function, optimized via cross-entropy. Simple and interpretable but limited for highly non-linear relationships.
XGBoost An optimized gradient boosting algorithm using second-order derivatives and regularization for efficiency and robustness. Fast and accurate, though often requires careful hyperparameter tuning.
CatBoost A gradient boosting method on decision trees that handles categorical features natively with ordered encodings. Strong performance on tabular data with minimal tuning.
MLP A Multi-Layer Perceptron applying three ReLU-activated fully connected layers, outputting class probabilities through a Softmax head.
DilatedCNN (1D) Stacks dilated 1D convolutions with normalization and dropout before a linear classification head producing class scores and Softmax probabilities from the final time step.
GraphCastGRU Encodes temporal signals with a GRU, feeds embeddings into GraphCast, and applies linear layers plus Softmax activation for graph-level class predictions.
LSTM Processes sequences with an LSTM block (optionally layer-normalized), applies dropout and ReLU-activated linear layers, and outputs class probabilities via a Softmax classifier.
GRU Uses a stacked GRU to summarize temporal dynamics, applies normalization and dropout, then maps the hidden state through linear layers followed by Softmax for classification outputs.
@@
    Silva et al. [16] introduce AUOC, a metric that jointly captures classification accuracy and ranking error while accounting for class imbalance and unobserved categories. It is computed by tracing paths along the confusion-matrix diagonal, with a Benefit term rewarding large correct entries and a Penalty term proportional to the distance from the diagonal. A perfect model will give an auoc of 0. AUOC is less penalizing when a fire event is predicted as a nearby but incorrect class (e.g., predicting class 0 instead of 1) and vice versa.
  3. 3.
    Intersection over Union (IoU), which measures how well the predicted risk aligns with actual risk when an event occurs. IoU is well-suited for multiclass wildfire prediction as it accounts for class uncertainty and preserves class ordinality—predicting class 1 instead of 4 is penalized less than predicting 0. With this metric, a false prediction is penalized proportionally to the class (0 to 4 is worst than 0 to 1 and vice versa).